Shutdown ellipsomter SNF 2006-10-09 11:25:56: microscope lamp is not working anymore

emyers at emyers at
Wed Oct 25 10:58:42 PDT 2006

Ted replaced microscope and ACT bulbs.  I adjusted stage height (which was off) and ran Uli's calibration wafer.  Results thickness 1045 vs. target of 1045, index of 1.472 vs target of 1.47.

More information about the ellipsomter-pcs mailing list